# Spatial-Temporal-Spectral Unified Modeling for Remote Sensing Dense Prediction

### v2

@@ Abstract
The proliferation of multi-source remote sensing data has propelled the development of deep learning for dense prediction, yet significant challenges in data and task unification persist. Current deep learning architectures for remote sensing are fundamentally rigid. They are engineered for fixed input-output configurations, restricting their adaptability to the heterogeneous spatial, temporal, and spectral dimensions inherent in real-world data. Furthermore, these models neglect the intrinsic correlations among semantic segmentation, binary change detection, and semantic change detection, necessitating the development of distinct models or task-specific decoders. This paradigm is also constrained to a predefined set of output semantic classes, where any change to the classes requires costly retraining. To overcome these limitations, we introduce the Spatial-Temporal-Spectral Unified Network (STSUN) for unified modeling. STSUN can adapt to input and output data with arbitrary spatial sizes, temporal lengths, and spectral bands by leveraging their metadata for a unified representation. Moreover, STSUN unifies disparate dense prediction tasks within a single architecture by conditioning the model on trainable task embeddings. Similarly, STSUN facilitates flexible prediction across any set of semantic categories by integrating trainable category embeddings as metadata. Extensive experiments on multiple datasets with diverse STS configurations in multiple scenarios demonstrate that a single STSUN model effectively adapts to heterogeneous inputs and outputs, unifying various dense prediction tasks and diverse semantic class predictions. The proposed approach consistently achieves state-of-the-art performance, highlighting its robustness and generalizability for complex remote sensing applications.
@@ I INTRODUCTION
With the continuous advancement of remote sensing technologies and the increasing diversity of data acquisition methods [1], the field of remote sensing has entered a phase of rapid development [2]. Massive and multi-source remote sensing data have been widely applied to various dense prediction tasks, playing a crucial role in applications such as urban expansion monitoring [3], land cover classification [4], disaster damage assessment [5], crop classification and yield estimation [6], and environmental pollution monitoring [7]. Remote sensing imagery exhibits high heterogeneity across three key dimensions including spatial, temporal and spectral dimensions, posing significant challenges for unified processing due to variations in image size, temporal length and spectral bands in practical applications [8].
Remote sensing dense prediction refers to tasks where the goal is to produce a pixel-wise prediction, assigning a specific label to every pixel in the input satellite imagery [9]. From the task perspective, dense prediction in remote sensing primarily involves three core categories: semantic segmentation, binary change detection, and semantic change detection. These tasks can be formally defined as follows: given a remote sensing image time series of shape $(T_{1},C_{1},H_{1},W_{1})$, the deep learning model is expected to produce a prediction of shape $(T_{2},C_{2},H_{2},W_{2})$, where $T_{1}$ and $T_{2}$ denote the temporal lengths of the input and output, $C_{1}$ is the number of input channels, $C_{2}$ is the number of output classes, and $(H_{1},W_{1})$ and $(H_{2},W_{2})$ denote the image sizes of the input and output, respectively. Typically, $H_{1}=H_{2}$ and $W_{1}=W_{2}$. In semantic segmentation, the model aims to classify land cover types at a per-pixel level [9], corresponding to $T_{2}=1$ and $C_{2}\geq 2$; in binary change detection, the model identifies whether changes occur between adjacent time points [10], corresponding to $T_{2}=T_{1}-1$ and $C_{2}=2$; in semantic change detection, the model extracts land cover information at each time step to analyze semantic differences between any two time points [11], corresponding to $T_{2}=T_{1}$ and $C_{2}\geq 2$. Therefore, the temporal dimension $T_{1},T_{2}$ is strongly correlated with the type of dense prediction tasks, and the output spectral dimension $C_{2}$ is strongly correlated with the set of predicted semantic categories.
@@
While deep learning methods have achieved significant progress in remote sensing dense prediction, yielding high-performing models across semantic segmentation, binary change detection, and semantic change detection tasks [12, 13, 14, 15], current architectures exhibit several limitations. 1. Fixed Configurations: These models are typically designed for fixed input-output configurations, defined by specific image sizes $((H1,W1),(H2,W2))$, temporal lengths $(T1,T2)$ and spectral bands $(C_{1},C_{2})$. This rigidity restricts their adaptability to the diverse and heterogeneous remote sensing datasets encountered in practical applications, such as diverse satellite sensor data with varying spectral bands or multi-temporal imagery sequences with varying temporal lengths. 2. Fixed Task: These models neglect the inherent correlations among semantic segmentation, binary change detection, and semantic change detection tasks. This oversight often mandates the development of distinct models or dedicated task-specific decoders for each task, hindering unified processing pipelines. 3. Fixed Category Set: Existing approaches are generally constrained to dense prediction with a fixed set of output categories. Consequently, any alteration to the target dense prediction schema, even a slight change, often result in substantial performance degradation or complete incompatibility, necessitating extensive retraining or fine-tuning, which incurs considerable computational and temporal overhead.
@@
To overcome these limitations, this study introduces the Spatial-Temporal-Spectral Unified Network (STSUN), as illustrated in Figure 1. STSUN offers a novel framework for unified representation and modeling of remote sensing data across diverse spatial, temporal and spectral dimensions, solving the above issues in the following ways. 1. Unified Representation: STSUN effectively addresses the variability in data characteristics by leveraging inherent metadata from each dimension to achieve a unified representation. This capability allows STSUN to seamlessly adapt to input and output data with arbitrary image sizes, temporal lengths, and spectral bands, thereby leveraging a wide range of data to build high-performance unified models for dense prediction. 2. Unified Tasks: STSUN unifies semantic segmentation, binary change detection, and semantic change detection tasks within a single framework, which is accomplished by using a predefined and trainable task embedding set. By incorporating a selected task embedding and output temporal length into the output temporal dimension as metadata, the model is explicitly instructed to execute on the exact dense prediction task and the necessary output temporal length. This approach eliminates the need for separate models or task-specific decoders, enabling effective multi-task joint modeling and leveraging extensive data from multiple dense prediction tasks to improve model performance. 3. Flexible Category Set: STSUN facilitates flexible prediction across an arbitrary set of semantic categories, which is accomplished by using a predefined and trainable category embedding set. The selected category subset is integrated into the output channel dimension as metadata, which guides the model to identify the specific classes needed for a given prediction. This enables the model to adapt to multiple subsets of prediction categories without fine-tuning, applying to diverse remote sensing scenarios with various semantic categories.
@@
    We design a flexible category prediction method, which utilizes trainable semantic embeddings to enable the model to perform dense prediction for any specified set of output classes, removing the constraint of a fixed category schema and the associated costs of retraining.
@@ II-A Data Dimensionality in Remote Sensing Dense Prediction
Spectral dimensionality is another source of major variation. The number of spectral bands can range from a single panchromatic band to a few multispectral bands (e.g., 4-bands in NAIP imagery, 13 bands in Sentinel-2 MSI [21]), to hundreds of narrow, contiguous bands in hyperspectral sensors like AVIRIS [28] or the upcoming EnMAP mission [29]. Each sensor captures information from different portions of the electromagnetic spectrum, with varying band central wavelengths and bandwidths. This spectral diversity allows for the discrimination of different materials and land cover types based on their unique spectral signatures [30]. However, it also means that models developed for one sensor may not be directly applicable to data from another sensor without strategies to handle the differing channel counts and spectral information content [31]. A series of models have been proposed to use hypernetworks to unify the inputs with different spectral bands, but ignored the unification of the output spectral bands [32, 33].
@@ II-B Task Unification in Remote Sensing Dense Prediction
Despite their promise, existing multi-task unification efforts suffer from critical limitations that hinder their scalability and effectiveness. The first issue is a pervasive dependence on paired data. Current models typically require that the training datasets are fully annotated for all constituent tasks, which means the exact same set of images must possess corresponding pixel-level labels for semantic segmentation, binary change, and semantic change. This stringent requirement drastically limits the pool of usable training data, as such comprehensively annotated, multi-task datasets are exceedingly rare and expensive to create. The second issue is the common architectural choice of using task-specific decoders. Many unified models employ a shared encoder to extract features but diverge into separate, specialized decoder heads for each task. This design can create information bottlenecks and prevent the model from fully exploiting the synergies and complementarities among the tasks at the deepest levels of feature decoding and synthesis.
@@ III-B Overview of the Spatial-Temporal-Spectral Unified Network
Figure 2: Overview of the STSUN architecture, comprising the Spectral-Spatial Unified Module for input and output alignment, the Basic blocks for feature extraction, and the Temporal Unified Module for temporal fusion and output adaptation.
@@
First, at the input stage, the spatial-spectral unified module (SSUM) leverages the input spatial metadata $M_{spa}^{in}$ and spectral metadata $M_{spe}^{in}$ for unified encoding. The temporal dimension, being intrinsically linked to the specific dense prediction task, necessitates a distinct processing approach from the spatial and spectral dimensions. If the temporal dimension were to be unified at this stage, it would be equivalent to a pixel-level fusion of multi-temporal remote sensing images. This would introduce significant interference from irrelevant information, thereby degrading the model’s performance on dense prediction tasks [13, 12]. Therefore, at this stage, the independence of the temporal dimension is preserved while the spatial and spectral dimensions are unified. In the spatial dimension, the spatial resolutions of the input images and the spatial position of each pixel are utilized as metadata for the input spatial dimension. This enables a uniform representation of data based on the similarity of spatial proximity, accommodating varied resolutions. Concurrently, in the channel dimension, the spectral wavelength of each spectral band is incorporated as metadata for the input channel dimension, facilitating a unified data representation grounded in spectral continuity.
Second, at the encoder stage, the encoder local global blocks are used to extract features from input data. As the temporal independence of remote sensing images is explicitly preserved, a shared-weight encoder then processes each temporal remote sensing image independently, extracting local and global features across distinct temporal instances within encoder local global blocks.
@@
Fourth, the decoder subsequently processes the fused multi-temporal remote sensing image features. Since the features from multiple temporal instances have been fused and the model has been instructed on the dense prediction task via the task embedding, the shared-weight decoder optimizes the features for the specific task by removing irrelevant information, progressively aligning them with the target output.
@@
Fifth, in the output stage, the SSUM utilizes the output spatial metadata $M_{spa}^{out}$ to restore the features to the original image dimensions and leverages the output spectral metadata $M_{spe}^{out}$ to guide the features toward making dense predictions for a specific subset of semantic categories. In the spatial dimension, the image size of the remote sensing images and the pixel position information are again leveraged as metadata for the output spatial dimension to accurately restore features to the original input image size. In the channel dimension, multiple predefined trainable category embeddings are established. From these, a subset of category embeddings is dynamically selected and integrated as metadata for the output channel dimension, explicitly indicating the set of categories the model is instructed to predict, thus enabling flexible class-agnostic prediction.
@@
Through the arrangement of spatial, temporal, and spectral dimensions across these five stages, and the design of metadata for each dimension, STSUN accommodates diverse input/output dimensional configurations and various subsets of prediction classes. Moreover, these designs enables the effective joint modeling of semantic segmentation, binary change detection, and semantic change detection. By uncovering the correlations and complementarities among these tasks, STSUN leverages their combined data to enhance its performance across multiple dense prediction challenges.
III-C Dimension Unified Module
Figure 3: Architecture of the proposed Dimension Unified Module.
To facilitate the mapping from the original data space to the unified feature space across different dimensions (spatial, temporal and spectral), and transform the unified features into output of appropriate shapes according to the specific task and category requirements, STSUN introduces the Dimension Unified Module (DUM), which serves as the fundamental implementation of both SSUM and TUM. DUM employs metadata from STS variable subsets and trainable embeddings to generate adaptive weights and biases for linear layers, thereby enabling adaptive and metadata-aware feature transformation. As illustrated in the Figure 3, given an input tensor $X\in\mathbb{R}^{D_{1}\times H\times W}$ and its metadata $M\in\mathbb{R}^{D_{1}}$, as well as trainable embeddings $E\in\mathbb{R}^{D_{1}}$ when performing on output temporal and spectral dimension, where $D_{1}$ can refer to spatial, temporal or spectral dimension and varies across different tasks and scenarios, and $M$ is an metadata of $\{M_{spa}^{in},M_{tem}^{in},M_{spe}^{in},M_{spa}^{out},M_{tem}^{out},M_{spe}^{out}\}$. DUM maps the varying input tensor to a flexible feature $Y\in\mathbb{R}^{L\times D_{2}}$ through a hypernetwork generation process. The detailed procedure is as follows:
  • Metadata Embedding: Metadata $M$ is subjected to tokenization through a linear layer and positional encoding. A learnable class token [CLS] is prepended to the token sequence $T\in\mathbb{R}^{(D_{1}+1)\times d}$, where $d$ denotes the embedding dimension.
  • Cross-Variable Relationship Modeling: The metadata token sequence is processed through multiple transformer blocks to capture the underlying relationships among metadata tokens.
  • Adaptive Parameter Generation: The [CLS] token is linearly projected to generate bias parameters $b\in\mathbb{R}^{D_{2}}$, while the remaining tokens are projected to produce the weight matrix $W\in\mathbb{R}^{D_{1}\times D_{2}}$. The resulting $W$ and $b$ form a linear layer that maps the input features with $D_{1}$ channels to output features with $D_{2}$ channels. The input $X$ of shape $(L,D_{1})$ can be processed through the generated linear layer, resulting in an output of shape $(L,D_{2})$. Specifically, when this module is applied to the output temporal and spectral dimensions, an additional learnable embedding $E$ is fused with $X$ prior to linear layer processing, which indicate the specific dense prediction task required for the features (corresponding to the output temporal dimension) or the specific set of semantic categories to be predicted (corresponding to the output spectral dimension).
By adaptively mapping STS variable subsets to a unified feature representation, DUM not only standardizes the feature shape but also preserves important relationships present in the original remote sensing data. Additionally, DUM can transform this unified feature representation into an output of a specific shape as required by the task, based on the dimensionality metadata of the output data. This enables STSUN to effectively support inputs and outputs with various dimensional settings, allowing it to generalize to diverse dense prediction tasks and class sets.

### v3

@@ Abstract
The proliferation of multi-source remote sensing data has propelled the development of deep learning for dense prediction, yet significant challenges in data and task unification persist. Current deep learning architectures for remote sensing are fundamentally rigid. They are engineered for fixed input-output configurations, restricting their adaptability to the heterogeneous spatial, temporal, and spectral dimensions inherent in real-world data. Furthermore, these models neglect the intrinsic correlations among semantic segmentation, binary change detection, and semantic change detection, necessitating the development of distinct models or task-specific decoders. This paradigm is also constrained to a predefined set of output semantic classes, where any change to the classes requires costly retraining. To overcome these limitations, we introduce the Spatial-Temporal-Spectral Unified Network (STSUN) for unified modeling. STSUN can adapt to input and output data with arbitrary spatial sizes, temporal lengths, and spectral bands by leveraging their metadata for a unified representation. Moreover, STSUN unifies disparate dense prediction tasks within a single architecture by conditioning the model on trainable task embeddings. Similarly, STSUN facilitates flexible prediction across multiple set of semantic categories by integrating trainable category embeddings as metadata. Extensive experiments on multiple datasets with diverse Spatial-Temporal-Spectral configurations in multiple scenarios demonstrate that a single STSUN model effectively adapts to heterogeneous inputs and outputs, unifying various dense prediction tasks and diverse semantic class predictions. The proposed approach consistently achieves state-of-the-art performance, highlighting its robustness and generalizability for complex remote sensing applications.
@@ I INTRODUCTION
With the continuous advancement of remote sensing technologies and the increasing diversity of data acquisition methods [1], the field of remote sensing has entered a phase of rapid development [2]. Massive and multi-source remote sensing data have been widely applied to various dense prediction tasks, playing a crucial role in applications such as urban expansion monitoring [3], land cover classification [4], disaster damage assessment [5], crop classification and yield estimation [6], and environmental pollution monitoring [7]. Remote sensing imagery and label exhibits high heterogeneity across three key dimensions including spatial, temporal and spectral dimensions in dense prediction tasks, posing significant challenges for unified processing due to variations in image size, temporal length and spectral bands of input and output in practical applications [8].
Remote sensing dense prediction refers to tasks where the goal is to produce a pixel-wise prediction, assigning a specific label to every pixel in the input satellite imagery [9]. Dense prediction in remote sensing primarily involves three core categories: semantic segmentation, binary change detection, and semantic change detection. These tasks can be formally defined as follows: given a remote sensing image time series of shape $(T_{1},C_{1},H_{1},W_{1})$, a dense prediction model is expected to produce a prediction of shape $(T_{2},C_{2},H_{2},W_{2})$, where $T_{1}$ and $T_{2}$ denote the temporal lengths of the input and output, $C_{1}$ is the number of input channels, $C_{2}$ is the number of output classes, and $(H_{1},W_{1})$ and $(H_{2},W_{2})$ denote the image sizes of the input and output, respectively. Typically, $H_{1}=H_{2}$ and $W_{1}=W_{2}$. From the task perspective, in semantic segmentation, the model aims to classify land cover types at a per-pixel level [9], corresponding to $T_{2}=1$ and $C_{2}\geq 2$; in binary change detection, the model identifies whether changes occur between adjacent time points [10], corresponding to $T_{2}=T_{1}-1$ and $C_{2}=2$; in semantic change detection, the model extracts land cover information at each time step to analyze semantic differences between adjacent two time points [11], corresponding to $T_{2}=T_{1}$ and $C_{2}\geq 2$. Therefore, the temporal dimension $T_{1},T_{2}$ is correlated with the type of dense prediction tasks, and the output spectral dimension $C_{2}$ is correlated with the set of predicted semantic categories. From the data perspective, the input’s spatial dimension is related to the geographical coverage and spatial resolution of remote sensing data; the input’s temporal dimension correlates with the temporal coverage and temporal resolution of remote sensing data; the input’s spectral dimension corresponds to the modality of remote sensing data. This underscores the importance of unified modeling of inputs and outputs in the spatial-temporal-spectral dimensions to achieve a unified approach for remote sensing dense prediction tasks.
@@
While deep learning methods have achieved significant progress in remote sensing dense prediction, yielding high-performing models across semantic segmentation, binary change detection, and semantic change detection tasks [12, 13, 14, 15], current architectures exhibit several limitations. 1. Fixed Configurations: These models are typically designed for fixed input-output configurations, defined by specific image sizes $((H_{1},W_{1}),(H_{2},W_{2}))$, temporal lengths $(T_{1},T_{2})$ and spectral bands $(C_{1},C_{2})$. This rigidity restricts their adaptability to the diverse and heterogeneous remote sensing datasets encountered in practical applications, such as diverse satellite sensor data with varying spectral bands or multi-temporal imagery sequences with varying temporal lengths. 2. Fixed Task: These models neglect the inherent correlations among semantic segmentation, binary change detection, and semantic change detection tasks. This oversight often mandates the development of distinct models or dedicated task-specific decoders for each task, hindering unified processing pipelines. 3. Fixed Category Set: Existing approaches are generally constrained to dense prediction with a fixed set of output categories. Consequently, any alteration to the target dense prediction schema, even a slight change, often result in substantial performance degradation or complete incompatibility, necessitating extensive retraining or fine-tuning, which incurs considerable computational and temporal overhead.
@@
To overcome these limitations, this study introduces the Spatial-Temporal-Spectral Unified Network (STSUN), as illustrated in Figure 1. STSUN offers a novel framework for unified representation and modeling of remote sensing data across diverse spatial, temporal and spectral dimensions, solving the above issues in the following ways. 1. Unified Representation: STSUN effectively addresses the variability in data characteristics by leveraging inherent metadata from each dimension to achieve a unified representation. This capability allows STSUN to seamlessly adapt to input and output data with arbitrary image sizes, temporal lengths, and spectral bands, thereby leveraging a wide range of data to build high-performance unified models for dense prediction. 2. Unified Tasks: STSUN unifies semantic segmentation, binary change detection, and semantic change detection tasks within a single framework, which is accomplished by using a predefined and trainable task embedding set. By incorporating a selected task embedding and output temporal length into the output temporal dimension as metadata, the model is explicitly instructed to execute on the exact dense prediction task and the necessary output temporal length. This approach eliminates the need for separate models or task-specific decoders, enabling effective multi-task joint modeling and leveraging extensive data from multiple dense prediction tasks to improve model performance. 3. Flexible Category Set: STSUN facilitates flexible prediction across multiple sets of semantic categories, which is accomplished by using a predefined and trainable category embedding set. The selected category subset is integrated into the output channel dimension as metadata, which guides the model to identify the specific classes needed for a given prediction. This enables the model to adapt to multiple subsets of prediction categories without fine-tuning, applying to diverse remote sensing scenarios with various semantic categories.
@@
    We design a flexible category prediction method, which utilizes trainable semantic embeddings to enable the model to perform dense prediction for multiple specified set of output classes, removing the constraint of a fixed category schema and the associated costs of retraining.
@@ II-A Data Dimensionality in Remote Sensing Dense Prediction
Spectral dimensionality is another source of major variation. The number of spectral bands can range from a single panchromatic band to a few multispectral bands (e.g., 4-bands in NAIP imagery, 13 bands in Sentinel-2 MSI [21]), to hundreds of narrow, contiguous bands in hyperspectral sensors like AVIRIS [28] or the upcoming EnMAP mission [29]. Each sensor captures information from different portions of the electromagnetic spectrum, with varying band central wavelengths and bandwidths. This spectral diversity allows for the discrimination of different materials and land cover types based on their unique spectral signatures [30]. However, it also means that models developed for one sensor may not be directly applicable to data from another sensor without strategies to handle the differing spectral bandss and spectral information content [31]. A series of models have been proposed to use hypernetworks to unify the inputs with different spectral bands, but ignored the unification of the output spectral bands [32, 33].
@@ II-B Task Unification in Remote Sensing Dense Prediction
Despite their superior performance, existing multi-task unification efforts suffer from critical limitations that hinder their scalability and effectiveness. The first issue is a pervasive dependence on paired data. Current models typically require that the training datasets are fully annotated for all constituent tasks, which means the exact same set of images must possess corresponding pixel-level labels for semantic segmentation, binary change, and semantic change. This stringent requirement drastically limits the pool of usable training data, as such comprehensively annotated, multi-task datasets are exceedingly rare and expensive to create. The second issue is the common architectural choice of using task-specific decoders. Many unified models employ a shared encoder to extract features but diverge into separate, specialized decoder heads for each task. This design can create information bottlenecks and prevent the model from fully exploiting the synergies and complementarities among the tasks at the deepest levels of feature decoding and synthesis.
@@ III-B Overview of the Spatial-Temporal-Spectral Unified Network
Figure 2: Overview of the STSUN architecture, comprising the Input Spatial-Spectral Unified Module for spatial and spectral unification of input, the Local Global Basic blocks for global and local feature extraction, the Temporal Unified Module for temporal unification of input and output, and the Output Spatial-Spectral Unified Module for spatial and spectral unification of output. The highlighted portions in the shapes on either side of each unification module indicate the dimension that the module unifies.
@@
First, at the input stage, the input spatial-spectral unified module (ISSUM) leverages the input spatial metadata $M_{spa}^{in}$ and spectral metadata $M_{spe}^{in}$ for unified encoding of spatial and spectral dimensions. The temporal dimension, being intrinsically linked to the specific dense prediction task, necessitates a distinct processing approach from the spatial and spectral dimensions. If the temporal dimension were to be unified at this stage, it would be equivalent to a pixel-level fusion of multi-temporal remote sensing images. This would introduce significant interference from irrelevant information, thereby degrading the model’s performance on dense prediction tasks [13, 12]. Therefore, at this stage, the independence of the temporal dimension is preserved while the spatial and spectral dimensions are unified. In the spatial dimension, the spatial resolutions of the input images and the spatial position of each pixel are utilized as metadata for the input spatial dimension. This enables a uniform representation of data based on the similarity of spatial proximity, accommodating varied resolutions. Concurrently, in the channel dimension, the spectral wavelength of each spectral band is incorporated as metadata for the input channel dimension, facilitating a unified data representation grounded in spectral continuity.
Second, at the encoder stage, the Encoder Local Global Blocks are used to extract features from input data. As the temporal independence of remote sensing images is explicitly preserved, a shared-weight encoder then processes each temporal remote sensing image independently, extracting local and global features across distinct temporal instances within Encoder Local Global Blocks. The overall structure of the Encoder Local Global Blocks is the same as the transformer block, except that the global attention is replaced with local-global window attention, as shown in Figure 4.
@@
Fourth, the decoder subsequently processes the fused multi-temporal remote sensing image features. Since the features from multiple temporal instances have been fused and the model has been instructed on the dense prediction task via the task embedding, the shared-weight decoder local-global blocks optimize the features for the specific task by removing irrelevant information, progressively aligning them with the target output. The Decoder Local Global Block are the same as Encoder Local Global Block.
@@
Fifth, in the output stage, the output spatial-spectral unified module (OSSUM) utilizes the output spatial metadata $M_{spa}^{out}$ to restore the features to the original image dimensions and leverages the output spectral metadata $M_{spe}^{out}$ to guide the features toward making dense predictions for a specific subset of semantic categories. In the spatial dimension, the spatial resolution of the remote sensing images and the patch position information are again leveraged as metadata for the output spatial dimension to accurately restore features to the original input image size. In the channel dimension, multiple predefined trainable category embeddings are established. From these, a subset of category embeddings is dynamically selected and integrated as metadata for the output channel dimension, explicitly indicating the set of categories the model is instructed to predict, thus enabling flexible class-agnostic prediction.
@@
Through the arrangement of spatial, temporal, and spectral dimensions across these five stages, and the design of metadata for each dimension, STSUN accommodates diverse input/output dimensional configurations and various subsets of prediction classes. Moreover, these designs enable the effective joint modeling of semantic segmentation, binary change detection, and semantic change detection. By uncovering the correlations and complementarities among these tasks, STSUN leverages the combined data of multiple tasks to enhance its performance across multiple dense prediction challenges.
III-C Spatial-Temporal-Spectral Unified Module
Figure 3: Illustration of the Spatial-Temporal-Spectral Unified Module. (a) Input Spatial-Spectral Unified Module. (b) Temporal Unified Module. (c) Output Spatial-Spectral Unified Module.
[…]
The core mechanism of STSUM involves using metadata from each dimension, along with optional trainable embeddings, to generate adaptive linear layers. This enables the transformation of variable input data into unified features or, conversely, the conversion of unified features into variable outputs tailored to specific requirements. The feature mapping mechanism is consistent across the modules and comprises two main branches: a hyper-network branch and a mapping network branch. In the hyper-network branch, metadata from each dimension is first tokenized using a linear layer and augmented with positional encodings. These tokens are then processed through several Transformer blocks to capture the latent relationships among them. Finally, another linear layer generates the parameters for the adaptive mapping network. In the mapping network branch, this dynamically generated network applies a linear transformation to the input features, thereby unifying the input data or generating the adaptive output. Although the underlying principle is similar, each module requires distinct and meticulous design considerations regarding feature shapes and parameter generation. This ensures that the modules can achieve generic and robust dimensional unification by processing different metadata for different mapping requirements across various dimensions.
III-C1 Input Spatial-Spectral Unified Module
[…]
III-C2 Temporal Unified Module
The TUM unifies the temporal dimension of the input and output data. It employs a similar mechanism with ISSUM to first map the variable input temporal length $T_{1}$ to a predefined, unified length $T$, and then map this unified length to a variable output length $T_{2}$ based on task requirements, as shown in Figure 3(b). For simplicity, let $C=C_{e}\cdot C_{a}$, $L=H\cdot W$. The TUM isolates the temporal dimension, obtaining a single data sample $X^{\prime}\in\mathbb{R}^{T_{1}\times L\times C}$, and reshapes it to $X^{\prime}\in\mathbb{R}^{(L\cdot C)\times T_{1}}$, thereby transposing the temporal dimension to the position of the channel dimension. This allows the TUM to unify the temporal dimension of the input data using a similar method of ISSUM. TUM utilizes the temporal position information of the input data as metadata, $M_{\text{tem}}^{\text{in}}\in\mathbb{R}^{T_{1}}$, to adaptively generate a linear layer composed of a bias parameter $b\in\mathbb{R}^{T}$ and a weight matrix $W\in\mathbb{R}^{T_{1}\times T}$. This mapping network transforms the temporally variable input $X^{\prime}\in\mathbb{R}^{(L\cdot C)\times T_{1}}$ into a unified representation $X^{\prime}\in\mathbb{R}^{(L\cdot C)\times T}$.
Subsequently, according to the requirements of the specific dense prediction task, the TUM needs to map the unified temporal length $T$ to a variable length $T_{2}$. Unlike the previous operation of mapping variable features to unified features, this process is reversed, which necessitates differences in the metadata and the generated parameters. Specifically, the TUM uses the output temporal length information and a selected task embedding as the output temporal metadata, $M_{\text{tem}}^{\text{out}}\in\mathbb{R}^{T_{2}}$. These are mapped through linear layers and then summed to form the metadata token sequence. The task embedding is a predefined, trainable embedding selected from a set, such as {semantic segmentation embedding, binary change detection embedding, semantic change detection embedding}, to specify the dense prediction task being performed. The metadata token sequence is then processed by a Transformer block to capture latent relationships between tokens. Following this, it passes through a linear layer to generate only the weight parameter $W\in\mathbb{R}^{T_{2}\times T}$, which is reshaped to $W\in\mathbb{R}^{T\times T_{2}}$ to serve as the weights of the mapping network. Finally, this bias-free mapping network transforms $X^{\prime}\in\mathbb{R}^{(L\cdot C)\times T}$ into $X^{\prime}\in\mathbb{R}^{(L\cdot C)\times T_{2}}$, which is then reshaped back to $X^{\prime}\in\mathbb{R}^{T_{2}\times L\times C}$, thus converting the unified temporal length $T$ into a variable length $T_{2}$ according to specific task demands. The bias parameter is not generated because the mapping network would require a variable bias $b\in\mathbb{R}^{T_{2}}$, which cannot be generated from a fixed CLS token through a fixed linear layer.
III-C3 Output Spatial-Spectral Unified Module
The OSSUM unifies the spatial and spectral dimensions of the output data. It first maps the unified image size $(H,W)$ to a variable size $(H_{2},W_{2})$ using spatial metadata, and then maps the unified spectral spectral bands $C$ to a variable count $C_{2}$ using spectral metadata, as illustrated in Figure 3(c). The dimensional mapping mechanism of the OSSUM is similar to that of the TUM for the output temporal dimension, with the main difference lying in the metadata. Specifically, for the spatial dimension, the hyper-network branch of the OSSUM uses the patch position and spatial resolution as metadata, $M_{\text{spa}}^{\text{out}}\in\mathbb{R}^{P_{h}\cdot P_{w}}$. After passing through linear layers and Transformer blocks, it generates a weight parameter $W\in\mathbb{R}^{C_{a}\times(P_{h}\cdot P_{w})}$ for the mapping network. Since the output spatial size is identical to the input spatial size, we have $M_{\text{spa}}^{\text{out}}=M_{\text{spa}}^{\text{in}}$. In the mapping network branch, the input feature $X^{\prime}\in\mathbb{R}^{(H\cdot W)\times(C_{e}\cdot C_{a})}$ is reshaped to $X^{\prime}\in\mathbb{R}^{(H\cdot W\cdot C_{e})\times C_{a}}$. It is then transformed by the mapping network to yield $X^{\prime}\in\mathbb{R}^{(H\cdot W\cdot C_{e})\times(P_{h}\cdot P_{w})}$. An ‘unpatchify‘ operation performs up-sampling, and a final reshape operation converts it to $X^{\prime}\in\mathbb{R}^{(H_{2}\cdot W_{2})\times C_{e}}$, producing a variable-sized output in the spatial dimension.
Similarly, in the spectral dimension, the OSSUM first selects a subset from a predefined set of semantic class embeddings based on the requirements of the prediction class set. This subset of embeddings indicates all the classes the model needs to predict. For instance, while multiple remote sensing scenes might require dense prediction for various land cover types, leading to a total semantic class embedding set like {Tree, Water, Soil, Road, Building, Background}, a specific building extraction task would only require selecting the subset {Building, Background}. This subset directs the model to classify land cover into these two categories, enabling effective building extraction. Therefore, the selected subset of semantic class embeddings serves as the output spectral metadata, $M_{\text{spe}}^{\text{out}}\in\mathbb{R}^{C_{2}\times C_{e}}$. Through a similar mechanism above, a weight parameter $W\in\mathbb{R}^{C_{e}\times C_{2}}$ is generated for the mapping network. In the mapping network branch, $X^{\prime}\in\mathbb{R}^{(H_{2}\cdot W_{2})\times C_{e}}$ undergoes a linear transformation by the mapping network to produce $X^{\prime}\in\mathbb{R}^{(H_{2}\cdot W_{2})\times C_{2}}$. Finally, a reshape operation yields the output $Y\in\mathbb{R}^{H_{2}\times W_{2}\times C_{2}}$, producing a variable output in the spectral dimension.